How many molecules of oxygen gas are present in a 0.343 g sample of oxygen gas? work included
1
Expert's answer
2020-11-23T07:34:37-0500

To get molecule we need to find mole first

Mole=mass/ molar mass

Mole=0.343/16=0.021

Mole=number of molecules/avogadro's constant

Molecules=mole x avogadro's constant

=0.021×6.02×10^23

=1.26×10^22 molecules